The Information Metric on Rational Maps

Michael K. Murray

Abstract

The information metric is a construction in statistics which can be used to define a (possibly degenerate) metric on various moduli spaces such as those of instantons and harmonic maps. This metric is shown to be nondegenerate for the space of harmonic maps of the two-sphere onto itself of any degree.